OpenStudy (anonymous):

lim as x approaches 0 for (x+sinx)/x

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x+\sin(x)}{x}=\lim_{x \rightarrow 0}(\frac{x}{x}+\frac{\sin(x)}{x})\]

myininaya (myininaya):

that is a major hint

myininaya (myininaya):

because sin(x)/x approaches what as x approaches 0?

myininaya (myininaya):

I'm asking you what the following is equal to: \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=?\]